# Mixture of Scope Experts at Test: Generalizing Deeper Graph Neural Networks with Shallow Variants

### v3

@@ A.2 GNNs with personalized scoping
Heuristics methods. The heuristics for personalized scoping originate from works that generalize personalized PageRank (PPR) to GNNs. PPNP [19] first introduces PPR as the final propagation matrix for decoupled GNNs. GBP [7] combines reverse push and random walks to approximate PPR propagation. NDLS [80] examines the smoothing effects in graph diffusion, noting that the level of smoothness should be node-specific. NDM [25] advances this by developing a unified diffusion kernel that extends PPR with the heat kernel and enables custom propagation steps following NDLS. To generalize personalized scoping to non-decoupled GNNs, ShaDow [78] proposes a design principle that decouples the scope from the model depth. For each node, a shallow scope is constructed using its neighboring nodes with the top-k PPR.
Learnable methods. However, these heuristics assume homophily and heavily rely on topological information, often falling short on heterophilous graphs. Recent research has explored parameterized techniques to address this issue. One line of work integrates personalized scoping into GNN architectures. We refer to these methods as soft personalized scoping since they typically learn node-dependent weights to control the scope. GeniePath [38] proposes a gated unit as the scope controller. GAMLP [82] uses the attention mechanism to enable personalized scoping on a decoupled GNN [16]. NW-GNN [64] further extends this to non-decoupled GNNs by proposing a node-wise architectural search. ACM-GCN [41] employs an alternative strategy by introducing additional identity channels beyond aggregation.
@@ A.3 Graph Mixture of Experts
Another line of work considers different depths of GNNs as different experts and develops a gating module to activate a small subset of experts for each input node. Policy-GCN [30] uses reinforcement techniques and takes the average accuracy of different depth models as rewards. GraphMoE [62] and DA-MoE [76] proposes a top-$K$ sparse gating technique for mixing multiple GNN experts. Additionally, Mowst [79] proposes separating rich self-features from informative neighborhoods by using a mixture of weak MLP and strong GNN experts. However, these methods require retraining the GNN models, which introduces significant overhead and may suffer from overfitting, potentially downgrading each model’s performance.
@@ Appendix B Proof
=$\underset{(a)}{=}\frac{\exp\left(\frac{({\mathbf{f}}_{u}-{\bm{\mu}}_{1,i}^{(L)})^{2}}{-2\sigma^{2}}\right)}{\exp\left(\frac{({\mathbf{f}}_{u}-{\bm{\mu}}_{1,i}^{(L)})^{2}}{-2\sigma^{2}}\right)+\exp\left(\frac{({\mathbf{f}}_{u}-{\bm{\mu}}_{2,i}^{(L)})^{2}}{-2\sigma^{2}}\right)},$
@@
$\Pr(y_{u}=c_{1}|{\mathbf{f}}_{u})-\Pr(y_{u}=c_{1}|{\mathbf{f}}_{v})=$ (8)
$\frac{\exp\left(\frac{({\mathbf{f}}_{u}-{\bm{\mu}}_{1,i}^{(L)})^{2}}{-2\sigma^{2}}\right)\exp\left(\frac{({\mathbf{f}}_{v}-{\bm{\mu}}_{2,j}^{(L)})^{2}}{-2\sigma^{2}}\right)-\exp\left(\frac{({\mathbf{f}}_{v}-{\bm{\mu}}_{1,j}^{(L)})^{2}}{-2\sigma^{2}}\right)\exp\left(\frac{({\mathbf{f}}_{u}-{\bm{\mu}}_{2,i}^{(L)})^{2}}{-2\sigma^{2}}\right)}{\left[\exp\left(\frac{({\mathbf{f}}_{u}-{\bm{\mu}}_{1,i}^{(L)})^{2}}{-2\sigma^{2}}\right)+\exp\left(\frac{({\mathbf{f}}_{u}-{\bm{\mu}}_{2,i}^{(L)})^{2}}{-2\sigma^{2}}\right)\right]\left[\exp\left(\frac{({\mathbf{f}}_{v}-{\bm{\mu}}_{1,j}^{(L)})^{2}}{-2\sigma^{2}}\right)+\exp\left(\frac{({\mathbf{f}}_{v}-{\bm{\mu}}_{2,j}^{(L)})^{2}}{-2\sigma^{2}}\right)\right]},$
@@
=$\underset{(a)}{=}(p_{i}-p_{j})\cdot\mathbb{E}_{\Pr(m)}\left[{\bm{\mu}}_{1,m}^{(L-1)}-{\bm{\mu}}_{2,m}^{(L-1)}\right]$
@@
Proof for Theorem 3.3. Theorem 1 in [46] provides a PAC-Bayes subgroup generalization bound for GNNs with one-hop aggregation. In our work, we extend their theorem to arbitrary hop to investigate how different scope sizes affect the generalization of different subgroups. Besides that, both theorem follows the same setting. To avoid replicates, we only provide the proof for the different parts and direct readers to the specific sections of [46] for the rest details and proofs.
Appendix F.4 in [46] provides complete proof for the proposed PAC-Bayes bound. There exists $0<\alpha<\frac{1}{4}$, we have
@@
where $D_{m,0}^{\gamma/2}(P;\lambda)$ is bounded by [46]’s Lemma 4.
Lemma 4 in [46] further depends on the bound of [46]’s Lemma 2 in Appendix E, which calls out the main difference compared with our theorem. By substituting [46]’s Lemma 2 with our Lemma B.1, we complete the proof for our Theorem 3.3
@@ C.2 Homophily Metrics
The homophily/heterophily metrics are widely used as graph properties to measure the probability of nodes with the same class connected to each other. This paper uses a node-wise homophily metric called node homophily [53]. Node homophily defines the fraction of neighbors that have the same class for each node:
@@
$H_{\text{node}}^{{\mathcal{G}}}=\frac{1}{|{\mathcal{V}}|}\sum_{v}{H_{\operatorname{node}}^{{\mathcal{G}}}}_{v}.$ (25)
There are also some other commonly used homophily metrics in the literature. Edge homophily [1, 84] measures the fraction of edges that connect nodes of the same class. Class homophily [37, 41] further addresses the sensitivity issue of edge homophily on graphs with imbalanced classes. For instance, genius dataset [37] has a majority class for roughly 80% of nodes, which can mislead edge homophily into classifying it as a homophilous graph ($H_{\operatorname{edge}}^{{\mathcal{G}}}=0.618$). In contrast, class homophily accurately identifies it as a heterophilous graph ($H_{\operatorname{class}}^{{\mathcal{G}}}=0.080$). However, class homophily does not satisfy some desired properties, such as asymptotic constant baseline and empty class tolerance. [55] proposed a variant called adjusted homophily to address this issue.
@@
As a pioneer work, graph convolutional network (GCN) [29] provides a layer-wise architecture that stacks feature propagation with linear transformation to approximate spectral graph convolutions. The $(l+1)-$th layer of GCN is defined as:
@@
where $\mathbf{h}_{u}^{(0)}={\bm{x}}_{u}$. Subsequent GNNs typically modify GCN in terms of aggregator and backbone. Some works develop more expressive aggregators [61, 14, 35]. For example, GAT [61] substituting the degree-normalized coefficients in GCN’s aggregator with learnable attention scores. Another line of works alters the backbone by decoupling aggregation from transformation [66, 19, 16, 82, 47] or adding skip connections between layers [69, 32, 8, 81]. For instance, SGC [66] eliminates the ${\text{ReLU}\left(\cdot\right)}$ function in the GCN layer to allow precomputing feature propagation. JKNet [69] directly maps the concatenated outputs of each layer $[\mathbf{h}_{v}^{(0)};\cdots;\mathbf{h}_{v}^{(L)}]$ to the prediction.
@@
Real-world graphs often exhibit a mix of homophilous and heterophilous patterns [37, 34, 46]. An ideal way to incorporate sufficient homophilous information is to set a personalized scope, allowing different nodes to have distinct scope sizes.
@@ Appendix D Supplementary Analysis for Deeper GNNs’ Failure
The depth dilemma. Increasing the GNN depth favorably increases the number of non-linear transformations and expands the scope of each node to exponentially incorporate more neighboring information. Existing studies also show that deeper GNNs are provably more expressive by using distance-based metrics [51, 26] and WL-based metrics [49, 10]. In practice, however, performance degradation is widely observed when going deep. Many novel architectures have been proposed to alleviate this issue, yet they achieve only marginal gains over shallow variants, which is undesirable given deeper GNNs’ superior expressive power and substantially higher computational costs. How to effectively leverage the depth remains an open question.
Limitation of existing solutions. While useful, current techniques for deeper GNNs face inherent trading-offs between these three types of errors. Residual connections (e.g., ResGCN) address the vanishing gradient issue but result in a larger generalization gap [13]. Learnable cross-layer gating modules (e.g., G2-GNN) achieve personalized scoping to alleviate over-smoothing, but they lead to an increase in training difficulties and the risk of overfitting. Meanwhile, GCNII attempts to prevent overfitting by mixing initial embeddings and adding identity mappings to weight matrices, but this comes at the cost of reduced expressivity.
@@ Appendix E Compare Expert-Gate Joint Training MoEs with Independent Train-then-Merge MoEs
Expert-gate joint training is widely used in current MoE methods for LLMs. Instead, Moscat employs a separate "train-then-merge" strategy which is optimized for GNN experts. Our work is motivated by the observation that directly transferring MoE designs from LLMs to GNNs faces inherent challenges. In LLMs, the gating module and experts are co-trained using Top-K sparse gating to reduce training and inference time while maintaining performance. However, for GNNs, no clear scaling law exists, and sparse gating plus more parameters do not necessarily translate into accuracy improvements. In our experiments with Moscat, we found that Top-K sparse gating (especially with small K) often results in an accuracy drop compared to dense gating. Recent work [9] also reveals that sparse gating can lead to performance losses on heterophilous datasets. Our independent-train-then-merge approach for GNN MoEs offers distinct advantages over expert-gate joint training:
@@
normalization { layer [3], batch [27], - }
@@
[…]
@@ F.1 Datasets
We evaluate Moscat on datasets with various homophily ratios. We used the filtered version [54] of Squirrel datasets, which removes all the duplicated nodes that share the same neighbors and labels. These duplicates exist widely in the train and test set and can cause data leakage. The amazon-ratings dataset is from [54]. For the Penn94, arxiv-year, genius and snap-patents datasets, we follow the same settings as [36] with 50%/25%/25% random splits for train/valid/test. We run 10 times on each of the 10 benchmark datasets. We follow the setup in [54], which does not convert the directed graphs to undirected graphs and does not use reverse edges since the outgoing neighbors might not be observed during real-world inference.
@@
Moscat. In this setting, we follow prior work [30, 63] to train the gating model on the original validation set. In our experiments, we sample 90% of the original validation set ${\mathcal{V}}_{\text{val}}$ as ${\mathcal{V}}_{\text{hold}}$ and use the remaining 10% for gating validation. In some cases we observe further gains by training the gating model on the full ${\mathcal{V}}_{\text{val}}$ and validating on ${\mathcal{V}}_{\text{train}}$. Thereby, we introduce a binary hyperparameter $\eta\in\{0,0.1\}$, where $\eta=0.1$ denotes the former split and $\eta=0$ the latter.
@@ F.3 Baselines
To evaluate the flexibility of Moscat, we select three classic homophilous GNNs (GCN [29], SGC [66], GAT [61]), and two state-of-the-art heterophilous GNNs (GCNII [8], ACM-GCN [41]) which cover comprehensive GNN architectural designs. We note that many scope mixing methods have assumptions about GNN architectures. For example, [18] can only be used on decoupled GNNs like SGC, while [41, 35, 71] are incompatible with GNNs using learnable aggregators like GAT. We further compare our methods with four models designed for node classification under heterophily: H2GCN [84], GPRGCN [11], FSGNN [47], MixHop [1]; three Graph Transformers: GraphGPS [56], SGFormer [67], Polynormer [15]; and three Graph MoEs: GMoE [62], Mowst [79], DA-MoE [76]; We also include skip-connection methods designed for deeper GNNs: Jumping Knowledge [69], GAMLP [82] and G2-GNN [58].
@@ F.4 Hyperparameters
We observe that normalization is important for GNNs, especially on larger datasets, and tuning the depth can provide substantial accuracy gains. With proper hyperparameter tuning, classic GNNs are strong baselines on heterophilous graphs, which aligns with recent works’ observation [54, 43]. When tuning the hyperparameters of the MLP in Moscat, we notice it is not sensitive to the number of transformation layers. Setting the number of layers to 3 works well in every setting. We also find that Moscat faces an accuracy drop when setting dropout or feature dropout to a value larger than 0. Therefore, we set the layers of MLP to 3 and dropout to zero for all settings, as shown in Table 10.
We further investigate tuning the Moscat-specific hyperparameters. Table 6 displays each hyperparameter’s explanation and search range. For the hidden dimension of Moscat, we notice that Moscat often works well when setting the dimension to 256. In smaller datasets such as Chameleon, we observe that a lower dimension is enough and enjoys better training. We limit the dimension for the snap-patents dataset to 128 due to the GPU memory constraint. Note that we set the upper bound of the Moscat maximum scope size $L_{\text{max}}$ to 6, rather than limiting it to a shallow 2 to 3 hop neighborhood [78] or expanding it to a global scope [34]. This is because (1) shallow scope does not contain sufficient homophily in heterophilous graphs [84], and (2) according to the six degrees of separation theorem, a 6-hop neighborhood is already large enough for many Wikipedia, citation, and social networks. Further increasing the scope size will only provide marginal improvement. As shown in Table 10, setting $L_{\text{max}}$ to 6 works best in most cases. $L_{\text{max}}$ can be used as a crucial parameter to trade off accuracy and overhead (see Appendix G.3 for more details).

### v4

@@ A.2 GNNs with personalized scoping
Heuristics methods. The heuristics for personalized scoping originate from works that generalize personalized PageRank (PPR) to GNNs. PPNP [appnp] first introduces PPR as the final propagation matrix for decoupled GNNs. GBP [gbp] combines reverse push and random walks to approximate PPR propagation. NDLS [ndls] examines the smoothing effects in graph diffusion, noting that the level of smoothness should be node-specific. NDM [nigcn] advances this by developing a unified diffusion kernel that extends PPR with the heat kernel and enables custom propagation steps following NDLS. To generalize personalized scoping to non-decoupled GNNs, ShaDow [shadow] proposes a design principle that decouples the scope from the model depth. For each node, a shallow scope is constructed using its neighboring nodes with the top-k PPR.
Learnable methods. However, these heuristics assume homophily and heavily rely on topological information, often falling short on heterophilous graphs. Recent research has explored parameterized techniques to address this issue. One line of work integrates personalized scoping into GNN architectures. We refer to these methods as soft personalized scoping since they typically learn node-dependent weights to control the scope. GeniePath [geniepath] proposes a gated unit as the scope controller. GAMLP [gamlp] uses the attention mechanism to enable personalized scoping on a decoupled GNN [sign]. NW-GNN [nw-gnn] further extends this to non-decoupled GNNs by proposing a node-wise architectural search. ACM-GCN [acmgcn] employs an alternative strategy by introducing additional identity channels beyond aggregation.
@@ A.3 Graph Mixture of Experts
Another line of work considers different depths of GNNs as different experts and develops a gating module to activate a small subset of experts for each input node. Policy-GCN [policy-gnn] uses reinforcement techniques and takes the average accuracy of different depth models as rewards. GraphMoE [graphmoe] and DA-MoE [damoe] proposes a top-$K$ sparse gating technique for mixing multiple GNN experts. Additionally, Mowst [mowst] proposes separating rich self-features from informative neighborhoods by using a mixture of weak MLP and strong GNN experts. However, these methods require retraining the GNN models, which introduces significant overhead and may suffer from overfitting, potentially downgrading each model’s performance.
@@ Appendix B Proof
$\underset{(a)}{=}\frac{\exp\left(\frac{({\mathbf{f}}_{u}-{\bm{\mu}}_{1,i}^{(L)})^{2}}{-2\sigma^{2}}\right)}{\exp\left(\frac{({\mathbf{f}}_{u}-{\bm{\mu}}_{1,i}^{(L)})^{2}}{-2\sigma^{2}}\right)+\exp\left(\frac{({\mathbf{f}}_{u}-{\bm{\mu}}_{2,i}^{(L)})^{2}}{-2\sigma^{2}}\right)},$
@@
$\hskip-17.07164pt\begin{aligned} &\Pr(y_{u}=c_{1}|{\mathbf{f}}_{u})-\Pr(y_{u}=c_{1}|{\mathbf{f}}_{v})=\\ &\frac{\exp\left(\frac{({\mathbf{f}}_{u}-{\bm{\mu}}_{1,i}^{(L)})^{2}}{-2\sigma^{2}}\right)\exp\left(\frac{({\mathbf{f}}_{v}-{\bm{\mu}}_{2,j}^{(L)})^{2}}{-2\sigma^{2}}\right)-\exp\left(\frac{({\mathbf{f}}_{v}-{\bm{\mu}}_{1,j}^{(L)})^{2}}{-2\sigma^{2}}\right)\exp\left(\frac{({\mathbf{f}}_{u}-{\bm{\mu}}_{2,i}^{(L)})^{2}}{-2\sigma^{2}}\right)}{\left[\exp\left(\frac{({\mathbf{f}}_{u}-{\bm{\mu}}_{1,i}^{(L)})^{2}}{-2\sigma^{2}}\right)+\exp\left(\frac{({\mathbf{f}}_{u}-{\bm{\mu}}_{2,i}^{(L)})^{2}}{-2\sigma^{2}}\right)\right]\left[\exp\left(\frac{({\mathbf{f}}_{v}-{\bm{\mu}}_{1,j}^{(L)})^{2}}{-2\sigma^{2}}\right)+\exp\left(\frac{({\mathbf{f}}_{v}-{\bm{\mu}}_{2,j}^{(L)})^{2}}{-2\sigma^{2}}\right)\right]},\end{aligned}$ (8)
@@
$\underset{(a)}{=}(p_{i}-p_{j})\cdot\mathbb{E}_{\Pr(m)}\left[{\bm{\mu}}_{1,m}^{(L-1)}-{\bm{\mu}}_{2,m}^{(L-1)}\right]$
@@
Proof for Theorem 3.3. Theorem 1 in [demystifying] provides a PAC-Bayes subgroup generalization bound for GNNs with one-hop aggregation. In our work, we extend their theorem to arbitrary hop to investigate how different scope sizes affect the generalization of different subgroups. Besides that, both theorem follows the same setting. To avoid replicates, we only provide the proof for the different parts and direct readers to the specific sections of [demystifying] for the rest details and proofs.
Appendix F.4 in [demystifying] provides complete proof for the proposed PAC-Bayes bound. There exists $0<\alpha<\frac{1}{4}$, we have
@@
where $D_{m,0}^{\gamma/2}(P;\lambda)$ is bounded by [demystifying]’s Lemma 4.
Lemma 4 in [demystifying] further depends on the bound of [demystifying]’s Lemma 2 in Appendix E, which calls out the main difference compared with our theorem. By substituting [demystifying]’s Lemma 2 with our Lemma B.1, we complete the proof for our Theorem 3.3
@@ C.2 Homophily Metrics
The homophily/heterophily metrics are widely used as graph properties to measure the probability of nodes with the same class connected to each other. This paper uses a node-wise homophily metric called node homophily [geomgcn]. Node homophily defines the fraction of neighbors that have the same class for each node:
@@
$H_{\text{node}}^{{\mathcal{G}}}=\frac{1}{|{\mathcal{V}}|}\sum_{v}\text{h}_{\text{node}}[v].$ (25)
There are also some other commonly used homophily metrics in the literature. Edge homophily [mixhop, h2gcn] measures the fraction of edges that connect nodes of the same class. Class homophily [linkx, acmgcn] further addresses the sensitivity issue of edge homophily on graphs with imbalanced classes. For instance, genius dataset [linkx] has a majority class for roughly 80% of nodes, which can mislead edge homophily into classifying it as a homophilous graph ($H_{\operatorname{edge}}^{{\mathcal{G}}}=0.618$). In contrast, class homophily accurately identifies it as a heterophilous graph ($H_{\operatorname{class}}^{{\mathcal{G}}}=0.080$). However, class homophily does not satisfy some desired properties, such as asymptotic constant baseline and empty class tolerance. [platonov2024characterizing] proposed a variant called adjusted homophily to address this issue.
@@
As a pioneer work, graph convolutional network (GCN) [gcn] provides a layer-wise architecture that stacks feature propagation with linear transformation to approximate spectral graph convolutions. The $(l+1)$-th layer of GCN is defined as:
@@
where $\mathbf{h}_{u}^{(0)}={\bm{x}}_{u}$. Subsequent GNNs typically modify GCN in terms of aggregator and backbone. Some works develop more expressive aggregators [gat, pna, atp]. For example, GAT [gat] substituting the degree-normalized coefficients in GCN’s aggregator with learnable attention scores. Another line of works alters the backbone by decoupling aggregation from transformation [sgc, appnp, sign, gamlp, fsgnn] or adding skip connections between layers [jknet, resgcn, gcnii, degradation]. For instance, SGC [sgc] eliminates the $\text{ReLU}\left(\cdot\right)$} function in the GCN layer to allow precomputing feature propagation. JKNet [jknet] directly maps the concatenated outputs of each layer $\mathbf{h}_{v}^{(0)};\cdots;\mathbf{h}_{v}^{(L)}]$ to the prediction.
@@
Real-world graphs often exhibit a mix of homophilous and heterophilous patterns [linkx, global, demystifying]. An ideal way to incorporate sufficient homophilous information is to set a personalized scope, allowing different nodes to have distinct scope sizes.
@@ Appendix D Supplementary Analysis for Deeper GNNs’ Failure
The depth dilemma. Increasing the GNN depth favorably increases the number of non-linear transformations and expands the scope of each node to exponentially incorporate more neighboring information. Existing studies also show that deeper GNNs are provably more expressive by using distance-based metrics [oono2019graph, huang2020tackling] and WL-based metrics [morris2019weisfeiler, chen2020can]. In practice, however, performance degradation is widely observed when going deep. Many novel architectures have been proposed to alleviate this issue, yet they achieve only marginal gains over shallow variants, which is undesirable given deeper GNNs’ superior expressive power and substantially higher computational costs. How to effectively leverage the depth remains an open question.
Limitation of existing solutions. While useful, current techniques for deeper GNNs face inherent trading-offs between these three types of errors. Residual connections (e.g., ResGCN) address the vanishing gradient issue but result in a larger generalization gap [provable]. Learnable cross-layer gating modules (e.g., G2-GNN) achieve personalized scoping to alleviate over-smoothing, but they lead to an increase in training difficulties and the risk of overfitting. Meanwhile, GCNII attempts to prevent overfitting by mixing initial embeddings and adding identity mappings to weight matrices, but this comes at the cost of reduced expressivity.
@@ Appendix E Compare Expert-Gate Joint Training MoEs with Independent Train-then-Merge MoEs
Expert-gate joint training is widely used in current MoE methods for LLMs. Instead, Moscat employs a separate "train-then-merge" strategy which is optimized for GNN experts. Our work is motivated by the observation that directly transferring MoE designs from LLMs to GNNs faces inherent challenges. In LLMs, the gating module and experts are co-trained using Top-K sparse gating to reduce training and inference time while maintaining performance. However, for GNNs, no clear scaling law exists, and sparse gating plus more parameters do not necessarily translate into accuracy improvements. In our experiments with Moscat, we found that Top-K sparse gating (especially with small K) often results in an accuracy drop compared to dense gating. Recent work [chen2025mixture] also reveals that sparse gating can lead to performance losses on heterophilous datasets. Our independent-train-then-merge approach for GNN MoEs offers distinct advantages over expert-gate joint training:
@@
normalization { layer [layer_norm], batch [batch_norm], - }
@@
Base GNN
for GNN-Moscat($\ast$)
GCNII MixHop GAT GAT ACMGCN MixHop GCNII GAT
@@ F.1 Datasets
We evaluate Moscat on datasets with various homophily ratios. We used the filtered version [critical] of Squirrel datasets, which removes all the duplicated nodes that share the same neighbors and labels. These duplicates exist widely in the train and test set and can cause data leakage. The amazon-ratings dataset is from [critical]. For the Penn94, arxiv-year, genius and snap-patents datasets, we follow the same settings as [lim2021large] with 50%/25%/25% random splits for train/valid/test. We run 10 times on each of the 10 benchmark datasets. We follow the setup in [critical], which does not convert the directed graphs to undirected graphs and does not use reverse edges since the outgoing neighbors might not be observed during real-world inference.
@@
Moscat. In this setting, we follow prior work [policy-gnn, wang2021confident] to train the gating model on the original validation set. In our experiments, we sample 90% of the original validation set ${\mathcal{V}}_{\text{val}}$ as ${\mathcal{V}}_{\text{hold}}$ and use the remaining 10% for gating validation. In some cases we observe further gains by training the gating model on the full ${\mathcal{V}}_{\text{val}}$ and validating on ${\mathcal{V}}_{\text{train}}$. Thereby, we introduce a binary hyperparameter $\eta\in\{0,0.1\}$, where $\eta=0.1$ denotes the former split and $\eta=0$ the latter.
@@ F.3 Baselines
To evaluate the flexibility of Moscat, we select three classic homophilous GNNs (GCN [gcn], SGC [sgc], GAT [gat]), and two state-of-the-art heterophilous GNNs (GCNII [gcnii], ACM-GCN [acmgcn]) which cover comprehensive GNN architectural designs. We note that many scope mixing methods have assumptions about GNN architectures. For example, [nai] can only be used on decoupled GNNs like SGC, while [acmgcn, atp, alt] are incompatible with GNNs using learnable aggregators like GAT. We further compare our methods with four models designed for node classification under heterophily: H2GCN [h2gcn], GPRGCN [gprgnn], FSGNN [fsgnn], MixHop [mixhop]; three Graph Transformers: GraphGPS [graphgps], SGFormer [sgformer], Polynormer [polynormer]; and three Graph MoEs: GMoE [graphmoe], Mowst [mowst], DA-MoE [damoe]; We also include skip-connection methods designed for deeper GNNs: Jumping Knowledge [jknet], GAMLP [gamlp] and G2-GNN [g2gnn].
@@ F.4 Hyperparameters
We observe that normalization is important for GNNs, especially on larger datasets, and tuning the depth can provide substantial accuracy gains. With proper hyperparameter tuning, classic GNNs are strong baselines on heterophilous graphs, which aligns with recent works’ observation [critical, nid]. When tuning the hyperparameters of the MLP in Moscat, we notice it is not sensitive to the number of transformation layers. Setting the number of layers to 3 works well in every setting. We also find that Moscat faces an accuracy drop when setting dropout or feature dropout to a value larger than 0. Therefore, we set the layers of MLP to 3 and dropout to zero for all settings, as shown in Table 10.
We further investigate tuning the Moscat-specific hyperparameters. Table 6 displays each hyperparameter’s explanation and search range. For the hidden dimension of Moscat, we notice that Moscat often works well when setting the dimension to 256. In smaller datasets such as Chameleon, we observe that a lower dimension is enough and enjoys better training. We limit the dimension for the snap-patents dataset to 128 due to the GPU memory constraint. Note that we set the upper bound of the Moscat maximum scope size $L_{\text{max}}$ to 6, rather than limiting it to a shallow 2 to 3 hop neighborhood [shadow] or expanding it to a global scope [global]. This is because (1) shallow scope does not contain sufficient homophily in heterophilous graphs [h2gcn], and (2) according to the six degrees of separation theorem, a 6-hop neighborhood is already large enough for many Wikipedia, citation, and social networks. Further increasing the scope size will only provide marginal improvement. As shown in Table 10, setting $L_{\text{max}}$ to 6 works best in most cases. $L_{\text{max}}$ can be used as a crucial parameter to trade off accuracy and overhead (see Appendix G.4 for more details).